# Unified Multi-task Learning for Voice-Based Detection of Diverse Clinical Conditions

### v1

@@ 1. Introduction
The human voice, intrinsically natural, non-invasive, and readily collectable, is increasingly recognized as a potent and low-cost digital biomarker for health (Fagherazzi et al., 2021; Ramanarayanan et al., 2022). Acoustic properties embedded within speech—such as pitch, prosody, speaking rate, and loudness—extend beyond conveying linguistic and emotional states, offering valuable indicators of an individual’s cognitive, neurological, respiratory, and vocal well-being (Luz et al., 2021). These insights underpin the potential of voice analysis as a scalable modality for the early screening and monitoring of diverse conditions, including Parkinson’s disease (PD), mild cognitive impairment (MCI), airway-related diseases, and various voice disorders (Tsanas et al., 2012).
@@
The clinical manifestations of these conditions often leave discernible imprints on vocal production. For instance, respiratory ailments like airway stenosis can lead to abnormal breath control and altered vocal tract coordination (Gelbard et al., 2016), while voice disorders such as vocal fold paralysis or structural lesions typically result in unstable vocal quality (Roy et al., 2007). Neurological or cognitive conditions like PD or MCI can impact prosody and fluency, often yielding slowed or monotonous speech patterns (Skodda et al., 2011; Mart´ınez-Nicolás et al., 2021). Traditional diagnostic workflows frequently necessitate in-person visits to healthcare facilities or the use of cumbersome physiological sensors. In contrast, speech-based analysis presents a significantly lower-burden, remote-friendly alternative, amenable to passive or active collection during everyday interactions, thereby easing patient engagement and facilitating longitudinal monitoring (Robin et al., 2020).
@@
The proliferation of mobile devices equipped with local processing capabilities has paved the way for deploying speech-based diagnostic systems in real-world, privacy-sensitive settings. Crucially, raw speech recordings, which may contain sensitive personal information, do not need to be transmitted. Instead, acoustic representations like Mel-frequency cepstral coefficients (MFCCs) or log-Mel spectrograms can be extracted locally on-device (Bäckström, 2023). This on-edge processing paradigm safeguards patient privacy, promotes ethical data handling, and enables the scalable deployment of voice-based health technologies across diverse healthcare ecosystems.
@@
In this work, we introduce a multitask disorder classification framework, MARVEL  (Multi-task Acoustic Representations for Voice-based Health Analysis), that operates exclusively on such derived speech features.An overview of the our framework is shown in Figure 1. Our model employs a dual-branch architecture with a shared acoustic backbone and task-specific binary classification heads, facilitating task-conditioned inference across a spectrum of voice-related disorders. This multitask learning (MTL) structure promotes beneficial feature sharing across related disease tasks, potentially enhancing generalization, especially for patients presenting with co-existing conditions. We train and rigorously evaluate our framework on the Bridge2AI-Voice v2.0 dataset (Bensoussan et al., 2025; Goldberger et al., 2000), a large-scale, ethically sourced repository of speech samples linked to diagnostic labels for neurological, respiratory, and vocal disorders. Notably, this dataset provides pre-extracted acoustic features and de-identified metadata, aligning perfectly with our privacy-conscious modeling approach that obviates the need for raw audio access. In addition, voice-based analysis lends itself to a wide range of real-world clinical applications. It can be integrated into remote monitoring systems to track patients with chronic or progressive conditions such as heart failure or neurodegenerative diseases. In triage settings, brief voice recordings may assist in early screening and prioritization of at-risk individuals. Moreover, such non-invasive and low-cost tools are particularly valuable in low-resource environments, where access to specialized clinicians or diagnostic infrastructure is limited. Together, these use cases underscore the practical value of speech as a scalable modality for population-level health assessment.
@@
To complement our MARVEL  framework, we further investigate the clinical relevance of established acoustic biomarkers by conducting an auxiliary analysis examining the correlation between the learned deep embeddings from MARVEL  and traditional handcrafted acoustic features. Previous studies show that handcrafted features alone can offer reasonable discrimination across certain disorder subtypes, motivating their use as a reference point for model interpretation (Low et al., 2020; Vásquez-Correa et al., 2019). This post hoc analysis reveals measurable similarities between the two, providing additional interpretability into what our model has implicitly learned and showing that it captures clinically relevant vocal patterns in a richer and more nuanced manner than can be described by predefined acoustic features. Moreover, the better performance of MARVEL  compared to models trained directly on handcrafted features suggests that our model leverages these acoustic cues while also learning more complex and task-specific representations.
@@ 2.1. Acoustic Approaches to Vocal Health Assessment
The utility of voice as a non-invasive modality for diagnosing a wide array of conditions, spanning neurological, respiratory, and voice-related disorders, has garnered significant research interest in recent years (Kim et al., 2023; Gómez-Vilda et al., 2013; Tong and Sataloff, 2022). Voice-based biomarkers are particularly compelling due to their inherent accessibility, low cost, and sensitivity to subtle alterations in motor and respiratory function. For instance, Parkinson’s disease (PD) frequently manifests through characteristic changes such as monotone prosody and imprecise articulation(Brin, 1992; Canter, 1963). Similarly, chronic respiratory diseases like Chronic Obstructive Pulmonary Disease (COPD) are often associated with distinct alterations in respiratory acoustics, potentially indicating deficits in breath control and upper airway coordination (Gross et al., 2009). Given the substantial global prevalence of such conditions—COPD alone affects over 300 million people globally (Vogelmeier et al., 2017), and neurological disorders like PD impact a rapidly growing aging population (Aarsland et al., 2021)—there is an urgent and unmet need for scalable, remote, and non-invasive screening and monitoring tools. Consequently, voice analysis presents a timely and clinically relevant avenue to support early detection and ongoing management across diverse disease domains.
@@
Initial explorations in this field predominantly relied on handcrafted acoustic features—such as jitter, shimmer, harmonic-to-noise ratio (HNR), and formants—coupled with traditional machine learning classifiers like Support Vector Machines (SVMs) and Gaussian Mixture Models (GMMs) (Anagnostopoulos et al., 2015; Verde et al., 2018). While these seminal studies established foundational links between vocal anomalies and clinical conditions, they often faced challenges in terms of robustness and generalizability across diverse datasets and recording conditions. The advent of deep learning has revolutionized the field, enabling automatic feature extraction directly from acoustic representations like MFCCs and spectrograms. Convolutional Neural Networks (CNNs) and recurrent architectures, such as Bidirectional Long Short-Term Memory networks (BiLSTMs), have proven effective in modeling the complex spatial and temporal patterns inherent in these representations (Peng et al., 2023; Alam et al., 2025). More recently, large-scale self-supervised learning (SSL) models, pre-trained on vast amounts of unlabeled audio data, such as wav2vec 2.0 (Baevski et al., 2020) and Whisper (Radford et al., 2023), have demonstrated remarkable performance in various speech processing tasks, including applications in medical audio analysis (Cai et al., 2024; Violeta et al., 2022).
@@
To leverage complementary information from different acoustic views or modalities, fusion-based architectures have also been explored. For example, multibranch CNN or Temporal Convolutional Network (TCN) models have been applied to tasks like pathological breath sound classification (Zhao et al., 2022; Kim et al., 2024). However, many prior works employing fusion tend to use homogeneous encoder structures (e.g., identical CNN backbones for different inputs), potentially limiting their capacity to leverage architecture-specific inductive biases suited for distinct feature types. Furthermore, a common limitation in speech-based disease detection is the assumption of single-label classification, which overlooks the clinical reality that patients often exhibit overlapping symptoms or co-morbidities across multiple systems (e.g., concurrent vocal and respiratory issues).
@@
Our work differentiates itself by proposing a structurally heterogeneous dual-branch fusion model. This model processes MFCCs and log-Mel spectrograms through distinct, specialized encoders—ResNet and EfficientNet, respectively. This architectural choice is informed by recent findings demonstrating the unique strengths of these backbones for specific clinical audio tasks: Yu et al. (Yin et al., 2024) highlighted ResNet’s efficacy in capturing temporal patterns for dysarthria severity modeling via its residual connections, while Kumar et al. (Kumar et al., 2024) showcased EfficientNet’s advantages in spatial modeling of log-Mel spectrograms for depression detection through compound scaling. By integrating these complementary encoder types within a unified framework, we aim to capture a richer, more comprehensive acoustic representation. Crucially, our model operates exclusively on these derived acoustic features, ensuring privacy-conscious deployment without reliance on raw audio, a critical consideration for clinical applications.
@@ 2.2. Multitask Learning in Clinical AI and Voice Diagnostics
Multitask learning (MTL) has emerged as a powerful paradigm for simultaneously modeling multiple clinically relevant targets, particularly when these tasks are correlated and can benefit from shared underlying feature representations (Caruana, 1997). In the broader healthcare domain, MTL has been successfully applied to diverse problems, including prediction tasks using electronic health records (EHRs), patient outcome modeling, and multimodal clinical reasoning (Niu et al., 2025; Ding et al., 2019). Such models typically employ shared encoder layers (backbone) followed by task-specific output heads, often leading to improvements in sample efficiency, model robustness, and generalization performance due to the regularizing effect of shared knowledge.
In the context of voice-based diagnostics, however, the majority of existing models remain confined to single-task setups (Orozco-Arroyave et al., 2016; Govindu and Palwe, 2023), often assuming that each speech recording corresponds to a singular disease label. This simplification does not adequately reflect real-world clinical scenarios where patients frequently present with symptoms spanning multiple conditions, such as co-occurring vocal and respiratory disorders. Single-task models may struggle to disentangle these complex, overlapping patterns. MTL offers a natural and effective solution by explicitly modeling shared vocal biomarkers across different tasks while simultaneously learning to retain task-specific nuances necessary for accurate discrimination.
While recent studies have begun to explore the application of MTL to pathological speech analysis, for instance, in jointly assessing voice quality and speech intelligibility (Joshy and Rajan, 2023), several gaps remain. Many of these approaches still rely on homogeneous input types (e.g., using only spectrograms or only MFCCs for all tasks) or restrict their modeling scope to a single family of disorders (e.g., only neurological conditions). Furthermore, the full potential of leveraging complementary information from diverse acoustic representations within an MTL framework has not been thoroughly explored.
@@ 3.2. Dataset and Input Representations
Our study utilizes a curated subset of the Bridge2AI-Voice v2.0 dataset (Bensoussan et al., 2025; Goldberger et al., 2000), a large-scale, ethically sourced voice health database. This dataset encompasses over 19,000 voice recordings from 442 unique participants, collected across five clinical institutions in North America. Each participant performed a standardized set of prompted speech tasks, and these recordings are linked with clinical diagnoses. Critically, to safeguard participant privacy, the publicly accessible version of the dataset exclusively contains derived acoustic representations—including log-Mel spectrograms, MFCCs, and a variety of statistical voice descriptors—with all raw audio data excluded.
@@ 3.3.1. Modality-Specific Encoders
  • Spectrogram Encoder ($f_{\text{spec}}$): An EfficientNet-B0 (Tan and Le, 2019) architecture is employed to process the log-Mel spectrograms ($\mathbf{X}^{\text{spec}}$). To align with the standard 3-channel input expected by the ImageNet-pretrained model, the single-channel spectrogram is passed through an initial 2D convolutional layer that projects it to 3 channels before entering the EfficientNet backbone.
  • MFCC Encoder ($f_{\text{mfcc}}$): A ResNet18 (He et al., 2016) architecture is used for the MFCC inputs ($\mathbf{X}^{\text{mfcc}}$). The first convolutional layer of the ResNet18 is modified to accept single-channel input, corresponding to the MFCC feature dimension over time.
@@ 4.1. Experimental Setup
Data Augmentation: To enhance model generalization, distinct data augmentation techniques were applied during training. For MFCC inputs, we appliep zero-mean Gaussian noise $\mathcal{N}(0,\sigma^{2})$, where $\sigma=0.01$. Further, we employed frequency masking, where 8 consecutive frequency bins were masked, and time masking, where 20 consecutive time frames were masked. For log-Mel spectrogram inputs, we utilized SpecAugment (Park et al., 2019), which involved frequency masking (masking 15% of the frequency bins) and time masking (masking 15% of the time frames), in addition to the time warping procedure detailed in Section 3.3. These augmentations were applied stochastically exclusively during training.
@@ 4.2.2. Task-Specific Performance Analysis
Subtype Task Whisper (Radford et al., 2023) Wav2vec-BERT (Barrault et al., 2023) M-CTC-T (Lugosch et al., 2022) MARVEL
@@ 4.2.3. Comparison with Self-Supervised Learning Baselines
To further contextualize the performance of our supervised MARVEL  framework, we compare it against contemporary self-supervised learning (SSL) models commonly used in speech processing. Table 3 presents a task-by-task AUROC comparison with Whisper-small (Radford et al., 2023), Wav2Vec2-BERT (Barrault et al., 2023), and M-CTC-T (Lugosch et al., 2022). For fair comparison, these SSL models were fine-tuned on our specific downstream tasks using the same training data splits and by their design only utilize spectrograms.

### v2

# Unified Acoustic Representations for Screening Neurological and Respiratory Pathologies from Voice

@@ 1. Introduction
The human voice, intrinsically natural, non-invasive, and readily collectable, is increasingly recognized as a potent and low-cost digital biomarker for health (fagherazzi2021voice; naserclaes2021speech). Acoustic properties embedded within speech—such as pitch, prosody, speaking rate, and loudness—extend beyond conveying linguistic and emotional states, offering valuable indicators of an individual’s cognitive, neurological, respiratory, and vocal well-being (luz2021detecting). These insights underpin the potential of voice analysis as a scalable modality for the early screening and monitoring of diverse conditions, including Parkinson’s disease (PD), mild cognitive impairment (MCI), airway-related diseases, and various voice disorders (tsanas2011novel).
@@
The clinical manifestations of these conditions often leave discernible imprints on vocal production. For instance, respiratory ailments like airway stenosis can lead to abnormal breath control and altered vocal tract coordination (gelbard2016disease), while voice disorders such as vocal fold paralysis or structural lesions typically result in unstable vocal quality (roy2007toward). Neurological or cognitive conditions like PD or MCI can impact prosody and fluency, often yielding slowed or monotonous speech patterns (skodda2011vocal; martinez2021ten). Traditional diagnostic workflows frequently necessitate in-person visits to healthcare facilities or the use of cumbersome physiological sensors. In contrast, speech-based analysis presents a significantly lower-burden, remote-friendly alternative, amenable to passive or active collection during everyday interactions, thereby easing patient engagement and facilitating longitudinal monitoring (robin2020evaluation).
@@
The proliferation of mobile devices equipped with local processing capabilities has paved the way for deploying speech-based diagnostic systems in real-world, privacy-sensitive settings. Crucially, raw speech recordings, which may contain sensitive personal information, do not need to be transmitted. Instead, acoustic representations like Mel-frequency cepstral coefficients (MFCCs) or log-Mel spectrograms can be extracted locally on-device (backstrom2023privacy). This on-edge processing paradigm safeguards patient privacy, promotes ethical data handling, and enables the scalable deployment of voice-based health technologies across diverse healthcare ecosystems.
@@
In this work, we introduce a multitask disorder classification framework, MARVEL  (Multi-task Acoustic Representations for Voice-based Health Analysis), that operates exclusively on such derived speech features.An overview of the our framework is shown in Figure 1. Our model employs a dual-branch architecture with a shared acoustic backbone and task-specific binary classification heads, facilitating task-conditioned inference across a spectrum of voice-related disorders. This multitask learning (MTL) structure promotes beneficial feature sharing across related disease tasks, potentially enhancing generalization, especially for patients presenting with co-existing conditions. We train and rigorously evaluate our framework on the Bridge2AI-Voice v2.0 dataset (bensoussan2025bridge2ai; goldberger2000physiobank), a large-scale, ethically sourced repository of speech samples linked to diagnostic labels for neurological, respiratory, and vocal disorders. Notably, this dataset provides pre-extracted acoustic features and de-identified metadata, aligning perfectly with our privacy-conscious modeling approach that obviates the need for raw audio access. In addition, voice-based analysis lends itself to a wide range of real-world clinical applications. It can be integrated into remote monitoring systems to track patients with chronic or progressive conditions such as heart failure or neurodegenerative diseases. In triage settings, brief voice recordings may assist in early screening and prioritization of at-risk individuals. Moreover, such non-invasive and low-cost tools are particularly valuable in low-resource environments, where access to specialized clinicians or diagnostic infrastructure is limited. Together, these use cases underscore the practical value of speech as a scalable modality for population-level health assessment.
@@
To complement our MARVEL  framework, we further investigate the clinical relevance of established acoustic biomarkers by conducting an auxiliary analysis examining the correlation between the learned deep embeddings from MARVEL  and traditional handcrafted acoustic features. Previous studies show that handcrafted features alone can offer reasonable discrimination across certain disorder subtypes, motivating their use as a reference point for model interpretation (low2020automated; vasquez2019convolutional). This post hoc analysis reveals measurable similarities between the two, providing additional interpretability into what our model has implicitly learned and showing that it captures clinically relevant vocal patterns in a richer and more nuanced manner than can be described by predefined acoustic features. Moreover, the better performance of MARVEL  compared to models trained directly on handcrafted features suggests that our model leverages these acoustic cues while also learning more complex and task-specific representations.
@@ 2.1. Acoustic Approaches to Vocal Health Assessment
The utility of voice as a non-invasive modality for diagnosing a wide array of conditions, spanning neurological, respiratory, and voice-related disorders, has garnered significant research interest in recent years (kim2023non; gomez2013characterizing; tong2022respiratory). Voice-based biomarkers are particularly compelling due to their inherent accessibility, low cost, and sensitivity to subtle alterations in motor and respiratory function. For instance, Parkinson’s disease (PD) frequently manifests through characteristic changes such as monotone prosody and imprecise articulation(brin1992movement; canter1963speech). Similarly, chronic respiratory diseases like Chronic Obstructive Pulmonary Disease (COPD) are often associated with distinct alterations in respiratory acoustics, potentially indicating deficits in breath control and upper airway coordination (gross2009coordination). Given the substantial global prevalence of such conditions—COPD alone affects over 300 million people globally (vogelmeier2017global), and neurological disorders like PD impact a rapidly growing aging population (aarsland2021parkinson)—there is an urgent and unmet need for scalable, remote, and non-invasive screening and monitoring tools. Consequently, voice analysis presents a timely and clinically relevant avenue to support early detection and ongoing management across diverse disease domains.
@@
Initial explorations in this field predominantly relied on handcrafted acoustic features—such as jitter, shimmer, harmonic-to-noise ratio (HNR), and formants—coupled with traditional machine learning classifiers like Support Vector Machines (SVMs) and Gaussian Mixture Models (GMMs) (anagnostopoulos2015features; verde2018voice). While these seminal studies established foundational links between vocal anomalies and clinical conditions, they often faced challenges in terms of robustness and generalizability across diverse datasets and recording conditions. The advent of deep learning has revolutionized the field, enabling automatic feature extraction directly from acoustic representations like MFCCs and spectrograms. Convolutional Neural Networks (CNNs) and recurrent architectures, such as Bidirectional Long Short-Term Memory networks (BiLSTMs), have proven effective in modeling the complex spatial and temporal patterns inherent in these representations (Peng2023; alam2025revamping). More recently, large-scale self-supervised learning (SSL) models, pre-trained on vast amounts of unlabeled audio data, such as wav2vec 2.0 (baevski2020wav2vec) and Whisper (radford2023robust), have demonstrated remarkable performance in various speech processing tasks, including applications in medical audio analysis (Cai2024; violeta2022investigating).
@@
To leverage complementary information from different acoustic views or modalities, fusion-based architectures have also been explored. For example, multibranch CNN or Temporal Convolutional Network (TCN) models have been applied to tasks like pathological breath sound classification (zhao2022automatic; kim2024real). However, many prior works employing fusion tend to use homogeneous encoder structures (e.g., identical CNN backbones for different inputs), potentially limiting their capacity to leverage architecture-specific inductive biases suited for distinct feature types. Furthermore, a common limitation in speech-based disease detection is the assumption of single-label classification, which overlooks the clinical reality that patients often exhibit overlapping symptoms or co-morbidities across multiple systems (e.g., concurrent vocal and respiratory issues).
@@
Our work differentiates itself by proposing a structurally heterogeneous dual-branch fusion model. This model processes MFCCs and log-Mel spectrograms through distinct, specialized encoders—ResNet and EfficientNet, respectively. This architectural choice is informed by recent findings demonstrating the unique strengths of these backbones for specific clinical audio tasks: Yu et al. (yin2024sarnet) highlighted ResNet’s efficacy in capturing temporal patterns for dysarthria severity modeling via its residual connections, while Kumar et al. (kumar2024depression) showcased EfficientNet’s advantages in spatial modeling of log-Mel spectrograms for depression detection through compound scaling. By integrating these complementary encoder types within a unified framework, we aim to capture a richer, more comprehensive acoustic representation. Crucially, our model operates exclusively on these derived acoustic features, ensuring privacy-conscious deployment without reliance on raw audio, a critical consideration for clinical applications.
@@ 2.2. Multitask Learning in Clinical AI and Voice Diagnostics
Multitask learning (MTL) has emerged as a powerful paradigm for simultaneously modeling multiple clinically relevant targets, particularly when these tasks are correlated and can benefit from shared underlying feature representations (caruana1997multitask). In the broader healthcare domain, MTL has been successfully applied to diverse problems, including prediction tasks using electronic health records (EHRs), patient outcome modeling, and multimodal clinical reasoning (niu2025medical; ding2019effectiveness). Such models typically employ shared encoder layers (backbone) followed by task-specific output heads, often leading to improvements in sample efficiency, model robustness, and generalization performance due to the regularizing effect of shared knowledge.
In the context of voice-based diagnostics, however, the majority of existing models remain confined to single-task setups (Orozco2016PD; govindu2023early), often assuming that each speech recording corresponds to a singular disease label. This simplification does not adequately reflect real-world clinical scenarios where patients frequently present with symptoms spanning multiple conditions, such as co-occurring vocal and respiratory disorders. Single-task models may struggle to disentangle these complex, overlapping patterns. MTL offers a natural and effective solution by explicitly modeling shared vocal biomarkers across different tasks while simultaneously learning to retain task-specific nuances necessary for accurate discrimination.
While recent studies have begun to explore the application of MTL to pathological speech analysis, for instance, in jointly assessing voice quality and speech intelligibility (joshy2023dysarthria), several gaps remain. Many of these approaches still rely on homogeneous input types (e.g., using only spectrograms or only MFCCs for all tasks) or restrict their modeling scope to a single family of disorders (e.g., only neurological conditions). Furthermore, the full potential of leveraging complementary information from diverse acoustic representations within an MTL framework has not been thoroughly explored.
@@ 3.2. Dataset and Input Representations
Our study utilizes a curated subset of the Bridge2AI-Voice v2.0 dataset (bensoussan2025bridge2ai; goldberger2000physiobank), a large-scale, ethically sourced voice health database. This dataset encompasses over 19,000 voice recordings from 442 unique participants, collected across five clinical institutions in North America. Each participant performed a standardized set of prompted speech tasks, and these recordings are linked with clinical diagnoses. Critically, to safeguard participant privacy, the publicly accessible version of the dataset exclusively contains derived acoustic representations—including log-Mel spectrograms, MFCCs, and a variety of statistical voice descriptors—with all raw audio data excluded.
@@ 3.3.1. Modality-Specific Encoders
  • Spectrogram Encoder ($f_{\text{spec}}$): An EfficientNet-B0 (tan2019efficientnet) architecture is employed to process the log-Mel spectrograms ($\mathbf{X}^{\text{spec}}$). To align with the standard 3-channel input expected by the ImageNet-pretrained model, the single-channel spectrogram is passed through an initial 2D convolutional layer that projects it to 3 channels before entering the EfficientNet backbone.
  • MFCC Encoder ($f_{\text{mfcc}}$): A ResNet18 (he2016deep) architecture is used for the MFCC inputs ($\mathbf{X}^{\text{mfcc}}$). The first convolutional layer of the ResNet18 is modified to accept single-channel input, corresponding to the MFCC feature dimension over time.
@@ 4.1. Experimental Setup
Data Augmentation: To enhance model generalization, distinct data augmentation techniques were applied during training. For MFCC inputs, we appliep zero-mean Gaussian noise $\mathcal{N}(0,\sigma^{2})$, where $\sigma=0.01$. Further, we employed frequency masking, where 8 consecutive frequency bins were masked, and time masking, where 20 consecutive time frames were masked. For log-Mel spectrogram inputs, we utilized SpecAugment (park2019specaugment), which involved frequency masking (masking 15% of the frequency bins) and time masking (masking 15% of the time frames), in addition to the time warping procedure detailed in Section 3.3. These augmentations were applied stochastically exclusively during training.
@@ 4.2.2. Task-Specific Performance Analysis
Subtype Task Whisper (radford2023robust) Wav2vec-BERT (barrault2023seamless) M-CTC-T (lugosch2022pseudo) MARVEL
@@ 4.2.3. Comparison with Self-Supervised Learning Baselines
To further contextualize the performance of our supervised MARVEL  framework, we compare it against contemporary self-supervised learning (SSL) models commonly used in speech processing. Table 3 presents a task-by-task AUROC comparison with Whisper-small (radford2023robust), Wav2Vec2-BERT (barrault2023seamless), and M-CTC-T (lugosch2022pseudo). For fair comparison, these SSL models were fine-tuned on our specific downstream tasks using the same training data splits and by their design only utilize spectrograms.